MATH 376 – Groundrules for Collaboration on Problem Sets and Labs – 2/15/2012

After grading the first problem set and the first labs (and after some general departmental discussions on these issues), I think we have to clarify some “do’s” and “don’t’s” regarding the practice of collaboration on the problem sets and labs for this course. The underlying goal of these rules is to ensure you are all putting in the individual effort needed to master the material for this course in addition to the group “brainstorming” that can be a fun and productive way to work on generating ideas. I have come to realize that there are a number of practices going on in many of our upper-level classes that really undermine this goal. After all, our problem sets are designed to be vehicles for learning; they are not annoyances to be “gotten out of the way” with as little time and effort as possible.

0. If you feel that working on the problem sets is taking too much time, you may omit copying out the problem statements. However, your solutions should still have enough explanation that you can go back later and understand what you were doing.

1. Do not divide up the problems on an assignment among members of the group and then compile the solutions for the group members to submit by copying work supplied by your group-mates (i.e. even with proper citation of who is in your working group). (Reason: This may seem like a good labor-saving device. But in practice, it means that if you have not really worked through some of the problems yourself, then you may not have really understood what is going on in them. You may also be turning in work that contains errors you have not caught because you have not thought and worked through the solution individually!)

2. Do include notes giving credit for office hours help you have received directly (individually or with others).

3. Do not delegate members from your group to come to office hours to ask about specific questions, and then compile help I gave to those individuals into solutions that are submitted by all members of the group. Do not credit office hours or incorporate ideas I shared there in your solutions if you have not been there yourself. In other words, collaboration on a problem must cease when some members of a group have been to office hours asking about a problem and others have not been there. (Reason: I try to make office hours as interactive as possible, so that I am pointing you in the right direction to solving a problem rather than showing you the whole solution myself. I try to get you to supply many or most of the key ideas and many technical details. If you copy work that relies on the interaction with me and other students without having been there, you have not gone through that process yourself and you don’t have any right to present that work as your own.)

4. If anyone asks you to do something that violates these groundrules, don’t comply with that request. If you have a question about whether something you want to do is OK or violates these groundrules, ask Prof. Little.

I have read and understood these rules and intend to follow them for the remainder of MATH 376:

Signed: